let triangleABC have vertices A(3,4),B(4,6) and C(-1,6).The distance b/w orthocentre of triangle and the origin is

Grade:12

2 Answers

Ajay Verma
10 years ago
ans:

triangle ABC is a right angle triangle.. so the ortho-center of the triangle will be the vertex where the right angle made..which is vertex A (3,4) ( you can figure it out by drawing this triangle geometrically or by checking slopes. )

so ortho center is A(3,4)
distance b/w A and O (origin) = ( 3^2 + 4^2 )1/2 ( by distance formula )
= 5


Thanks and Regards,
Ajay verma,
askIITians faculty,
IIT HYDERABAD
Ajay Verma
10 years ago
ans:

triangle ABC is a right angle triangle.. so the ortho-center of the triangle will be the vertex where the right angle made..which is vertex A (3,4) ( you can figure it out by drawing this triangle geometrically or by checking slopes. )

so ortho center is A(3,4)
distance b/w A and O (origin) = ( 3^2 + 4^2 )1/2( by distance formula )
= 5

Thanks and Regards,
Ajay verma,
askIITians faculty,
IIT HYDERABAD